The equation of straight line passes through a point(4, 3)  and making intercepts on coordinate axes whose sum is -1

Grade:11

2 Answers

Saurabh Koranglekar
5 years ago
(x/a) + (y/b) = 1 be the equation of line

(4/a) + (3/b) = 1

also a+b = -1

two equation two variables

you can solve from here revert if problem still persists
Arun
5 years ago
 
Dear student
 
Let yhe kine is 
 
x/a + y/b = 1
 
Now we are given that a+ b = – 1
&
 
4/a + 3/b = 1
 
4b + 3 (- 1 – b) = (-1 – b) b
 
it is quadratic in b 
 
Now you can find b then a
 
hence you can also find the equation of line.